# Quantum Acoustics Demystifies the Strange Metals

### v1

@@ Abstract
Phonons have long been thought to be incapable of explaining key phenomena in strange metals, including linear-in-T Planckian resistivity from high to very low temperatures. We argue that these conclusions were based on static, perturbative approaches that overlooked essential time-dependent and nonperturbative electron-lattice physics. In fact “phonons” are not the best target for discussion, just like “photons” are not the best way to think about Maxwell’s equations. Quantum optics connects photons and electromagnetism, as developed 60 years ago by Glauber and others. We have been developing the parallel world of quantum acoustics. Far from being only of academic interest, the new tools are rapidly exposing the secrets of the strange metals, revealing strong vibronic (vibration-electronic) interactions playing a crucial role forming polarons and charge density waves, linear-in-$T$ resistivity at the Planckian rate over thousands of degrees, resolution of the Drude peak infrared anomaly, and the absence of a $T^{4}$ low-temperature resistivity rise in 2D systems, and of a Mott-Ioffe-Regel resistivity saturation. We derive Planckian transport, polarons, CDWs, and pseudogaps from the Fröhlich model. The “new physics” has been hiding in this model all along, in the right parameter regime, if it is treated nonperturbatively. In the course of this work we have uncovered the generalization of Anderson localization to dynamic media: a universal Planckian diffusion emerges, a “ghost” of Anderson localization. Planckian diffusion is clearly defined and is more fundamental than the popular but elusive, model dependent concept of “Planckian speed limit”.
@@ I Introduction
In the 1950s and 60s, the Hanbury-Brown Twiss experiment made it imperative to unify the particle-like photon picture of light with the wave-like electromagnetic field obeying Maxwell’s equations. Thus was born the new paradigm of quantum optics, where Glauber coherent states became the vehicle to travel from photons to electromagnetic waves.
@@
The opportunity to develop solid state theory of lattice vibration along parallel lines has been largely neglected. The phonon plays the role of the photon, and acoustic waves play the role of electromagnetic waves.
@@
The connection goes all the way back to Schrödinger, who worked out the number state-coherent state connection in 1926, Fig. 1 [schr]. The former is natural to the time-independent Schrödinger equation, and the latter needs the time-dependent Schrödinger equation.
@@
We have recently been developing this paradign for electron in a lattice, and have found it to be remarkably powerful for both physical insight and computations. We call it quantum acoustics. We do not need to introduce a new physical model, but rather we develop the Fröhlich model in a perfectly valid but neglected coherent state representation. New physics is found to emerge, as well-grounded as it is in quantum optics.
@@
The traditional solid state paradigm is one of particles, i.e. phonons. In special contexts, involving transducers for example, acoustic waves are familiar, but the unification of phonons and acoustic waves has been lacking. It is natural that the coherent state is again the right tool to unify the wave and particle pictures, this time in solid state physics. Far from being of only academic interest, quantum acoustics has made it possible to understand several of the high-temperature superconductor, above-$T_{c}$ “strange metal” mysteries, within a single theoretical context and within perhaps the most well established model of electron-vibration inteactions (Fröhlich). This paper reviews that progress, and points the way to likely future developments.
@@
Earlier, we put the time-dependent, real space “wave-on-wave,” or WoW code described below through its paces regarding ordinary metals [heller22]. It gave essentially exact agreement with the Bloch-Grüneisen, perturbative + Boltzmann transport theory, including a $T^{4}$ ($T^{5}$ in 3D) low temperature rise of resistivity at low $T$ [heller22]. Neither perturbation theory nor Boltzmann transport was used. Now we are applying quantum acoustics to the high $T_{c}$ materials in their strange “normal” (non-superconducting) state, and it is yielding a cornucopia of results. The WoW approach, including the lattice interacting with the electrons and vise-versa is nonperturbative and coherence preserving; both traits are needed to understand the strange metals.
Still, the community hoping for new electron-electron strogly correlated phyics has a refuge, for the moment, from these developments. We already think we can eventessentaill ually explain all the above Tc experimantal results, but we cannot yet claim to have pinned down the very low temperature linearity of resistivity. We do have palusible scenarios but this leads us into speculaiton, which mist of work avoids.
Inseed the results presented below are at a high level of confidence, so much so that any add-on theory of strongly coupled electron fluid must be just that: an add-on, that coexists with out quantum acoustics results.
@@
“Strongly correlated” behavior is often cited as a hallmark of the strange metals. In our perspective, there is a strong coupling that emerges within the venerable Fröhlich Hamiltonian here, and is vibronic in origin: Electrons become strongly coupled with each other because they are each strongly coupled to the same lattice and its deformations due to thermal vibrations and the presence of the electron back-action. Even near 0 Kelvin, given a deformation potential constant of 10 or 20 eV, strong coupling emerges in the strange metals (but not for the ordinary metals).
@@
The basis for this approach can be traced back to the semiclassical (regarding the lattice) polaron theory of Landau and Pekar [landau1948effective], as amplified by Leopold et.al.  [leopold2022landau]. In analogy with an electron in a classical electromagnetic field, Landau-Pekar takes a classical lattice approach as opposed to a quantized, number-state, phonon-based approach.
@@
Another important antecedent is Peierls distortion of a lattice. This very idea shifts us quietly into a coherent state representation of the lattice, because it is extremely inconvenient to think of the distortions in terms of phonon occupations. Sound waves in a crystal coming from a transducer also necessarily slip us into the coherent state representation, or at least a classical lattice picture.
@@
The Fröhlich Hamiltonian, with its centerpiece electron-phonon interaction term, generates strongly coupled, emergent behavior, if it is allowed to remain always “on,” acting coherently to unlimited order. The quantum acoustics paradigm can implement the Fröhlich Hamiltonian nicely, but it should be noted that the quantum acoustics framework is much more general than the Fröhlich model. That is, much higher order models can also be accomodated withing quantum mscoustics.
@@
The term “quantum acoustics” is not new, but it has typically meant controlled, engineered, and quantum-coherent manipulation of sound waves at the single-phonon or few-phonon level. Initially the term appeared in the quantum engineering / quantum information domain: coupling phonons (mechanical vibrations) with superconducting qubits, i.e. wave/phonon control at the quantum level. So, quantum acoustics has so far remained wedded to the phonon, i.e., a particle-like picture. In contrast, quantum optics encompasses both photons, which are particles, and waves, i.e., fields, as described in Maxwell’s equations. These disparate pictures are unified through the coherent state representation.
@@
There have been reasons for suspecting that lattice vibration could not be the root cause of several strange metal (high-$T_{c}$ material) mysteries, including the persistence of resistivity not only linear in T down to essentially 0 Kelvin, but also at the Planckian “speed limit” slope, $1/\tau=k_{B}T/(m^{*}D)$ This slope, if one is permitted to use Drude-like theory, corresponds to a Planckian relaxation time $\tau=\hbar/k_{B}T$. Thermal phonons acting alone in a flat medium unmarred by disturbances such as CDW or polarons are incompetent to cause such scattering at very low temperatures. In addition, it has been hard to imagine how the phonons could impose d-wave symmetry on the pseudogaps (we will see how this works through nesting vector influenced CDW, in the coherent state representation, below).
There were other arguments against a primary role for phonons (or better, lattice vibrations). The lack of isotope effects seemed to argue that the vibrations could not be involved, for example. However, there are problems with this conclusion. As discussed below, it may be incorrect to assume the deformation potential landscape is flat for an electron at low $T$. Lattice effects of the sort that lead to Peierls displacements and Kohn anomalies may be fluctuations, if not permanent features, of the landscape, even at very low T, and are examples of strong electron-lattice interaction.
As we shall show below, Planckian diffusion with diffusion constant $D=\hbar/m^{*}$ is ubiquitous, and is the successor to Anderson localization when the random medium is actively evolving. This fact is key to the strange metal regime. A strongly correlated electron-lattice chaotic “soup” emerges even from the Fröhlich model, and does not have to be postulated, as it is in the SYK model [kitaev2015, rosenhaus2019].
@@
Our quantum acoustic realization of the Fröhlich model (the wave-on-wave or WoW mean field approach, described in section II.2 ) derives its parameters directly from the best experimental estimates of the strange metals,. So far, we have hovered only near optimal doping. Within WoW, Planckian diffusion, spontaneous polaron formation, charge density waves, pseudogap behavior with d-wave symmetry, correct Drude peak displacement spectroscopy, and by-pass of the Mott-Ioffe-Regel limit all emerge naturally. This is making what can be called a strong case for lattice motion coupled to electrons as the root cause of the strange metal mysteries. Because the question was previously evaluated under a perturbative mindset, lattice vibrations have not been given a fair trial, until now.
@@ I.1 Tradition
The coherent state representation of lattice vibrations (leading to waves and fields) has received very little attention compared to the formally equivalent number state representation. In fact, coherent states have been discouraged, in spite of their extreme success and importance right next door in quantum optics, where the classical limit fields are governed by by Maxwell’s equations.
@@
We strongly disagree with “no new physics” claim. With such logic, there is “no new physics” in the time-dependent Schrödinger equation compared to the time-independent Schrödinger equation! That is only true in an extremely narrow sense, For most of us, “new physics” will mean the discovery of previously unknown and emergent phenomena not found before. Such discoveries are highly dependent on the starting representation, since exact solutions will never be available.
@@ II.1 Prior use of the term “quantum acoustics”
Before and concurrent with our work, the term “quantum acoustics” has not received wide use but has again usually meant a few quanta (phonons) such as “Listening For New Physics With Quantum Acoustics,” about a phonon-qubit swap device [linehan_listening_2024]. Or suface acoustic waves [9880558]. This is a bit different than the parallel context of “quantum optics”, which although often studied near the single photon limit, is also perfectly comfortable with classical electromagnetic fields. We are not suggesting a change in terminology anywhere, except we are broadening the term quantum acoustics to include everything from individual phonons, to classical limit acoustical waves. In this way it comports with quantum optics.
@@ II.2 Wave-on-wave method
Fröhlich Hamiltian
More than 70 years later, we have extended the program begun by Bardeen and Schockley [hellerkim], and Fröhlich, whose Hamiltonian is seen second-quantized form [Many-Particle], equation 1. It was always known it contained the potential for correctly describing very strong electron-lattice interactions, since Frölich’s first papers introduced and derived polarons with it [frohlich1954electrons].
@@
Within the present deformation potential framework, an electron undergoes continuous quasielastic, coherence-preserving deflection while roaming through the slowly altering (but Anderson localization destroying) potential landscape. The hills and valleys act like a smobile defect field, even in a perfect crystal, with the key twist that the defects are constantly morphing from one shape to another.
@@
Figure 6: Wave-on-Wave approach. Nonperturbative and coherent dynamics of charge carriers and thermal lattice vibrations. A snapshot of a carrier wave packet (the msgnitude is shown on top) coherently propagating under a spatially continuous, dynamic disorder field (bottom) formed by acoustic lattice deformations. A charge carrier quasielastically scatters (similar to impurity scattering) in this disordered landscape.
@@ III Two key findings with WoW
These findings comport with and enhance Hartnoll’s 2015 conjectures[Hartnoll2015] regarding diffusive transport at the Planckian rate $\tau=\hbar/k_{B}T$ in strange metals. The engine is indeed diffusion, but the fundamental element is Planckian diffusion $D=\hbar/m^{*}$ independent of temperature, see next section IV.
@@ V Planckian diffusion: The ghost of Anderson Localization
The discovery that Planckian diffusion rises from Anderson localization broken by motion of an otherwise localizing medium [ghost, Aydin24] has very wide implications, within and beyond the strange metals. To re-state: Quantum particles living in a random medium undergoing dynamic change, a medium that would Anderson localize the particles if it were stationary, will instead result in their diffusion with a constant $D\sim\hbar/m^{*},$ independent of temperature, independent of the coupling strength to the lattice. There will be little dependence on the isotopes present, or rate of mixing of the medium, or even the localization length in the absence of disturbances. This trend toward a kind of universality neutralizes many of the objections raised over the possible role of the lattice modes in strange metal behavior. The only system specific dependence is to the effective mass.
This Planckian diffusion is appropriately called a ghost[ghost], generalizing Anderson localization to moving media. It is more ubiquitous than Anderson localization, in the sense that the multitude of the localization scenarios is multiplied by the multitude of the ways of breaking them by activity of the medium. A huge variety of scenarios lead to approximately the same quantum diffusion constant.
@@
As one example, in the strange metal regime of the high $T_{c}$ superconductors, we have found numerically that diffusion of an electron wavepacket in the active deformation potential field lives within a quantum Planckian upper and lower bound, $D=\alpha\hbar/m^{*}$, where $m^{*}$ is the effective mass, and $1/2\leq\alpha\leq 2,$ independent of temperature or electron-phonon coupling strength [PNASpr]. Using our WoW code, we watch an electron diffuse at almost the same rate after temperature and coupling are changed by orders of magnitude.
As with the related Planckian speed limit, discussed next, we call the diffusion Planckian if $D$ is within a factor of 2 of $\hbar/m$. These upper and lower limits can be violated, so we call it a “ubiquitous” phenomenon. In our paper, “Planckian Diffusion: The ghost of Anderson Localization” [ghost], we demonstrate that once Anderson localization is destroyed by the real-time motion of a random field, an itinerant “ghost” emerges in its place: Planckian diffusion, characterized by $D\approx\hbar/m^{*}$. For this to apply, the medium must move quickly enough for the carrier to be transported non-adiabatically, an easily satisfied condition in most random media. In short, Planckian diffusion governs the strange metal phase: it is the leading cause and key symptom of strange metal behavior.

### v2

@@ Abstract
Phonons have long been thought to be incapable of explaining key phenomena in strange metals, including linear-in-T Planckian resistivity from high to very low temperatures. We argue that these conclusions were based on static, perturbative approaches that overlooked essential time-dependent and nonperturbative electron-lattice physics. In fact “phonons” are not the best target for discussion, just like “photons” are not the best way to think about Maxwell’s equations. Quantum optics connects photons and electromagnetism, as developed 60 years ago by Glauber and others. We have developed a parallel world of quantum acoustics. Far from being only of academic interest, the new tools are rapidly exposing the secrets of the strange metals, revealing strong vibronic (vibration-electronic) interactions playing a crucial role forming polarons and charge density waves, linear-in-$T$ resistivity at the Planckian rate over thousands of degrees, resolution of the Drude peak infrared anomaly, and the absence of a $T^{4}$ low-temperature resistivity rise in 2D systems, and of a Mott-Ioffe-Regel resistivity saturation. We derive Planckian transport, polarons, CDWs, and pseudogaps from the Fröhlich model. The “new physics” has been hiding in this model all along, in the right parameter regime, if it is treated nonperturbatively. In the course of this work we have uncovered the generalization of Anderson localization to dynamic media: a universal Planckian diffusion emerges, a “ghost” of Anderson localization. Planckian diffusion is clearly defined and is more fundamental than the popular but elusive, model dependent concept of “Planckian speed limit”.
@@ I Introduction
In the 1950s and 1960s, the Hanbury-Brown Twiss experiment made it imperative to unify the particle-like photon picture of light with the wave-like electromagnetic field obeying Maxwell’s equations. Thus was born the new paradigm of quantum optics, where Glauber coherent states became the vehicle to travel from photons to electromagnetic waves.
@@
The opportunity to develop solid state theory of lattice vibration along parallel lines has been largely neglected. The phonon plays the role of the photon, and the acoustic waves play the role of the electromagnetic waves.
@@
The connection goes back to Schrödinger, who worked out the number state-coherent state connection in 1926, Fig. 1 [schr]. The former is natural to the time-independent Schrödinger equation, and the latter needs the time-dependent Schrödinger equation.
@@
We have recently been developing this paradigm for electrons in a lattice and have found it to be remarkably powerful for both physical insight and computations. We call it quantum acoustics. We do not need to introduce a new physical model, but rather we develop the Fröhlich model in a perfectly valid but neglected coherent state representation. New physics emerges, as well-grounded as it is in quantum optics.
@@
The traditional solid state paradigm is one of particles, i.e. phonons. In special contexts, involving transducers, for example, acoustic waves are familiar, but the unification of phonons and acoustic waves has been lacking. It is natural that the coherent state is again the right tool to unify the wave and particle pictures, this time in solid state physics. Far from being of only academic interest, quantum acoustics has made it possible to understand several of the high-temperature superconductor, above-$T_{c}$ “strange metal” mysteries, within a single theoretical context and within perhaps the most well established model of electron-vibration interactions (Fröhlich). This paper reviews that progress and points the way to likely future developments.
@@
Earlier, we put the time-dependent real space “wave-on-wave,” or WoW code described below through its paces regarding ordinary metals [heller22]. It gave essentially exact agreement with the Bloch-Grüneisen, perturbative + Boltzmann transport theory, including a $T^{4}$ ($T^{5}$ in 3D) low temperature rise of resistivity at low $T$ [heller22]. Neither perturbation theory nor Boltzmann transport was used. Now we are applying quantum acoustics to the high $T_{c}$ materials in their strange “normal” (non-superconducting) state, and it is yielding a cornucopia of results. The WoW approach, including the lattice interacting with the electrons and vise-versa is nonperturbative and coherence preserving; both traits are needed to understand the strange metals.
We have already explained or believe we will be able to explain most of the above experimental results above Tc, but we cannot yet claim to have pinned down the very low temperature linearity of resistivity. We have plausible lattice-electron scenarios, but this leads us into speculation, which we mostly avoid here.
The veil is lifting on many of the strange metal mysteries and controversies, after nearly 40 years. It has become crystal clear that electron-lattice interactions are essential to understand the strange metal physics.
@@
“Strongly correlated” behavior is often cited as a hallmark of strange metals. In our perspective, there is a strong coupling that emerges within the venerable Fröhlich Hamiltonian, and is vibronic in origin: Electrons become strongly coupled indirectly with each other because they are each strongly coupled to the same lattice and its deformations due to thermal vibrations and the presence of the electron back-action. Even near 0 Kelvin, given a deformation potential constant of 10 or 20 eV, strong electron-lattice coupling emerges in strange metals (but not for ordinary metals).
@@
The basis for this approach can be traced back to the semiclassical (with respect to the lattice) polaron theory of Landau and Pekar [landau1948effective], as amplified by Leopold et al.  [leopold2022landau]. In analogy with an electron in a classical electromagnetic field, Landau-Pekar takes a classical lattice approach as opposed to a quantized, number-state, phonon-based approach.
@@
Another important antecedent is Peierls distortion of a lattice. This very idea shifts us quietly into a coherent state representation of the lattice because it is extremely inconvenient to think of the distortions in terms of phonon occupations. Sound waves in a crystal coming from a transducer also necessarily slip us into the coherent state representation, or at least a classical lattice picture.
@@
The Fröhlich Hamiltonian, with its centerpiece electron-phonon interaction term, generates strongly coupled emergent behavior, if it is allowed to remain always “on,” acting coherently to unlimited order. The quantum acoustics paradigm can implement the Fröhlich Hamiltonian nicely, but it should be noted that the quantum acoustics framework is much more general than the Fröhlich model. That is, much higher order models can also be accommodated with quantum acoustics.
@@
The term “quantum acoustics” is not new, but it has typically meant controlled, engineered, and quantum-coherent manipulation of sound waves at the single-phonon or few-phonon level. Initially, the term appeared in the quantum engineering / quantum information domain: coupling phonons (mechanical vibrations) with superconducting qubits, i.e., wave/phonon control at the quantum level. So far, quantum acoustics has remained wedded to the phonon, i.e., a particle-like picture. In contrast, quantum optics encompasses both photons, which are particles, and waves, i.e., fields, as described in Maxwell’s equations. These disparate pictures are unified through the coherent state representation.
@@
There have been reasons for suspecting that lattice vibration could not be the root cause of several strange metal (high-$T_{c}$ material) mysteries, including the persistence of resistivity not only linear in T down to essentially 0 Kelvin, but also at the Planckian “speed limit” slope, $1/\tau=k_{B}T/(m^{*}D).$ This slope, if one is permitted to use Drude-like theory, corresponds to a Planckian relaxation time $\tau=\hbar/k_{B}T$. Thermal phonons acting alone in a flat medium unmarred by disturbances such as CDW or polarons are incompetent to cause such scattering at very low temperatures. In addition, it has been hard to imagine how the phonons could impose d-wave symmetry on the pseudogaps (we will see how this works through nesting vector influenced CDW, in the coherent state representation, below).
There were other arguments against a primary role for phonons (or better, lattice vibrations). The lack of isotope effects seemed to argue that vibrations could not be involved, for example. However, there are problems with this conclusion. As discussed below, it may be incorrect to assume the deformation potential landscape is flat for an electron at low $T$. Lattice effects of the sort that lead to Peierls displacements and Kohn anomalies may be fluctuations, if not permanent features, of the landscape, even at very low T, and are examples of strong electron-lattice interaction.
As we shall show in the following, Planckian diffusion with diffusion constant $D=\hbar/m^{*}$ is ubiquitous and is the successor to Anderson localization when the random medium is actively evolving. This fact is key to the strange metal regime. A strongly correlated electron-lattice chaotic “soup” emerges even from the Fröhlich model, and does not have to be postulated, as it is in the SYK model [kitaev2015, rosenhaus2019].
@@
Our quantum acoustic realization of the Fröhlich model (the wave-on-wave or WoW mean field approach, described in section II.2 ) derives its parameters directly from the best experimental estimates of the strange metals,. So far, we have hovered only near optimal doping. Within WoW, Planckian diffusion, spontaneous polaron formation, charge density waves, pseudogap behavior with d-wave symmetry, correct Drude peak displacement spectroscopy, and by-pass of the Mott-Ioffe-Regel limit all emerge naturally. This is making a strong case for lattice motion coupled to electrons as the root cause of the strange metal mysteries. Because the question was previously evaluated under a perturbative mindset, lattice vibrations have not been given a fair trial, until now.
@@ I.1 Tradition
The coherent state representation of lattice vibrations (leading to waves and fields) has received very little attention compared to the formally equivalent number state representation. In fact, coherent states have been discouraged, in spite of their extreme success and importance right next door in quantum optics, where the classical limit fields are governed by Maxwell’s equations.
@@
We strongly disagree with “no new physics” claim. With such logic, there is “no new physics” in the time-dependent Schrödinger equation compared to the time-independent Schrödinger equation! That is only true in an extremely narrow sense. For most of us, “new physics” will mean the discovery of previously unknown and emergent phenomena not found before. Such discoveries are highly dependent on the starting representation since exact solutions will never be available.
@@ II.1 Prior use of the term “quantum acoustics”
Before and concurrent with our work, the term “quantum acoustics” has not received wide use but has again usually meant a few quanta (phonons) such as “Listening For New Physics With Quantum Acoustics,” about a phonon-qubit swap device [linehan_listening_2024]. Or surface acoustic waves [9880558]. This is a bit different than the parallel context of “quantum optics”, which although often studied near the single photon limit, is also perfectly comfortable with classical electromagnetic fields. We are not suggesting a change in terminology anywhere, except we are broadening the term quantum acoustics to include everything from individual phonons, to classical limit acoustical waves. In this way it comports with quantum optics.
@@ II.2 Wave-on-wave method
Fröhlich Hamiltonian
More than 70 years later, we have extended the program begun by Bardeen and Shockley [hellerkim], and Fröhlich, whose Hamiltonian is seen second-quantized form [Many-Particle], equation 1. It was always known it contained the potential for correctly describing very strong electron-lattice interactions, since Frölich’s first papers introduced and derived polarons with it [frohlich1954electrons].
@@
Within the present deformation potential framework, an electron undergoes continuous quasielastic, coherence-preserving deflection while roaming through the slowly altering (but Anderson localization destroying) potential landscape. The hills and valleys act like a mobile defect field, even in a perfect crystal, with the key twist that the defects are constantly morphing from one shape to another.
@@
Figure 6: Wave-on-Wave approach. Nonperturbative and coherent dynamics of charge carriers and thermal lattice vibrations. A snapshot of a carrier wave packet (the magnitude is shown on top) coherently propagating under a spatially continuous, dynamic disorder field (bottom) formed by acoustic lattice deformations. A charge carrier quasielastically scatters (similar to impurity scattering) in this disordered landscape.
@@ III Two key findings with WoW
These findings comport with and enhance Hartnoll’s 2015 conjectures[Hartnoll2015], and contributions by Nussinov and Chakrabarty[Nussinov_ann.phys_443_168970_2022, NUSSINOV2022168970] regarding diffusive transport at the Planckian rate $\tau=\hbar/k_{B}T$ in strange metals. The engine is indeed diffusion, but the fundamental element is Planckian diffusion $D=\hbar/m^{*}$ independent of temperature, see next section IV.
@@ V Planckian diffusion: The ghost of Anderson Localization
In our paper, “Planckian Diffusion: The ghost of Anderson Localization” [ghost], we demonstrate that once Anderson localization is destroyed by the real-time motion of a random field, an itinerant “ghost” emerges in its place: Planckian diffusion, characterized by $D\approx\hbar/m^{*}$. For this to apply, the medium must move quickly enough for the carrier to be transported non-adiabatically, an easily satisfied condition in most random media. In short, Planckian diffusion governs the strange metal phase: it is the leading cause and key symptom of strange metal behavior. This discovery has wide implications, within and beyond the strange metals. To re-state: Quantum particles living in a random medium undergoing dynamic change, in a medium that would Anderson localize the particles if it were stationary, will instead result in their diffusion with a constant $D\sim\hbar/m^{*},$ independent of temperature, independent of the coupling strength to the lattice. There will be little dependence on the isotopes present, or rate of mixing of the medium, or even the localization length in the absence of disturbances. This trend toward a kind of universality neutralizes many of the objections raised over the possible role of the lattice modes in strange metal behavior. The only system specific dependence is to the effective mass.
This Planckian diffusion is appropriately called a ghost[ghost], generalizing Anderson localization to moving media. It is more ubiquitous than Anderson localization, in the sense that the multitude of the localization scenarios is multiplied by the multitude of the ways of breaking them by activity of the medium. A huge variety of scenarios thus lead to approximately the same quantum diffusion constant.
@@
In the strange metal regime of the high $T_{c}$ superconductors, we have found numerically that diffusion of an electron wavepacket in the active deformation potential field lives within a quantum Planckian upper and lower bound, $D=\alpha\hbar/m^{*}$, where $m^{*}$ is the effective mass, and $1/2\leq\alpha\leq 2,$ independent of temperature or electron-phonon coupling strength [PNASpr]. Using our WoW code, we watch an electron diffuse at almost the same rate after the temperature and coupling changes by orders of magnitude.
As with the related Planckian speed limit $\tau\sim\hbar/k_{B}T$, we call the diffusion Planckian if $D$ is within a factor of 2 of $\hbar/m^{*}$. These upper and lower limits can be violated, so we call it a “ubiquitous” phenomenon.
The closely related Planckian speed limit has received much more attention. See, e.g., Refs. [Zaanen_Sci.post_6_061_2019, Hartnoll_rev.mod.phys_94_041002_2022, Mousatov_nat.phys_16_579_2020, Zhang_pnas_116_19869_2019, Lucas_phys.rev.lett_122_216601_2019, Gleis_phys.rev.lett_134_106501_2025]). This idea rests in a compelling but somewhat ad hoc way on the time-energy uncertainty relation, $\Delta t\Delta E\sim\hbar$, assuming the characteristic energy scale $\Delta E$ to be associated with thermal $k_{B}T$. Although exactly what physical process suffers the corresponding time scale $\Delta t\sim\tau$ is sometimes vague, it can be connected to momentum relaxation [Hartnoll_rev.mod.phys_94_041002_2022]. Nussinov and Chakrabarty have provided a more concise perspective[Nussinov_ann.phys_443_168970_2022, NUSSINOV2022168970].
Planckian diffusion $D=\hbar/m^{*}$ encompasses the Planckian speed limit $\tau\sim\hbar/k_{B}T$. It is more general in the sense that it applies to systems not possessing a temperature or any equilibrium, and it doesn’t require a model for what $\tau$ represents. The speed limit form for thermal systems is reached from $D=\hbar/m^{*}$ by an Einstein relation.